OpenStudy (supercalifragalisticexspeaalli):

can someone help me plzzz?? i will medal&fan anyone who helps

OpenStudy (_malice_):

Mentally or Physically

OpenStudy (supercalifragalisticexspeaalli):

Question 1 Is the expression x3•x3•x3 equivalent to x3•3•3? Why or why not? Explain your reasoning. Question 2 Rewrite in simplest radical form 1 x −3 6 . Show each step of your process. Question 3 Rewrite in simplest rational exponent form √x • 4√x. Show each step of your process. Question 4 Rewrite in simplest radical form x 5 6 x 1 6 . Show each step of your process. Question 5 Which of the following expressions are equivalent? Justify your reasoning. 4√x3 1 x−1 10√x5•x4•x2 x 1 3 •x 1 3 •x 1 3

OpenStudy (supercalifragalisticexspeaalli):

@Addy @agent0smith @AaronAndyson @mjdennis @_Malice_ @Natasha18

OpenStudy (supercalifragalisticexspeaalli):

i really need some help

OpenStudy (mjdennis):

I'll help with the first one. You need to repost the others, one question per post please, so everyone can help, everyone can get credit, and no one gets overwhelmed.

OpenStudy (supercalifragalisticexspeaalli):

ok np

rishavraj (rishavraj):

\[a^x \times a^y \times a^z = a^{x+y+z}\]

OpenStudy (mjdennis):

Question 1 Is the expression x3•x3•x3 equivalent to x3•3•3? Why or why not? Explain your reasoning. Your notation is poor, it could mean several things. Do you mean"\[x ^{3} * x ^{3} * x ^{3}\] and \[x ^{3*3*3}\] If so, you nbeed to type in exponents as "x^3"

OpenStudy (supercalifragalisticexspeaalli):

i think the first one isn't equivalent because it say's x^3*x^3*x^3*x^3 but the other one is x^3*3*3

OpenStudy (supercalifragalisticexspeaalli):

yea n sry bout that

OpenStudy (mathstudent55):

Is this the first question?